OpenStudy (anonymous):

How do i simplify : ln e^7 and e^3lnx

OpenStudy (anonymous):

ln e^7:ln and e cancel leaving only 7

OpenStudy (anonymous):

I think it's x^3 instead of 3x

OpenStudy (anonymous):

no i dont think so

OpenStudy (anonymous):

I meant the answer is x^3

OpenStudy (anonymous):

for e^3lnx

OpenStudy (anonymous):

Remember the exponential function and natural logarithm are inverses. \[e^{3 \ln x} = e^{\ln x^3} = x^3 \]

OpenStudy (anonymous):

what the natural log of 1

OpenStudy (anonymous):

0